OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of (f + g)(3)? −4 4 −32 8

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf f(x) = x^2 − 13 }\) \(\Large\color{blue}{ \bf f(g) = x+5 }\) \(\Large\color{blue}{ \bf f+g=(x^2-13)+(x+5)=x^2+x-8 }\)